What do "the product of two numbers is equal to 14" mean?..
What do "2 times a number equals that number plus 5" mean?..
What do "half of a number is 8 less than a second number" mean?
What do " 4 more than 3 times a number is equal to 6 more than 2 times the number" mean?
What do "a number decreased by 5 and then doubled is 4" mean?

Answer by checkley77(12569)

What do "the product of two numbers is equal to 14" mean?..
XY=14
What do "2 times a number equals that number plus 5" mean?..
2X=X+5
What do "half of a number is 8 less than a second number" mean?
X/2=Y-8
What do " 4 more than 3 times a number is equal to 6 more than 2 times the number" mean?
3X+4=2X+6
What do "a number decreased by 5 and then doubled is 4" mean?
(X-5)2=4